OpenStudy (anonymous):

Solve ln(5x+5)=2 round to the nearest thousand??

OpenStudy (anonymous):

Raise both sides to \(e\): \[e^{\ln(5x+5)} = e^2\]Do you know what happens when you have \(e^{\ln(x)}\)?

OpenStudy (anonymous):

Nice name @Richard_Feynman and explanation :)

OpenStudy (anonymous):

@Richard_Feynman no

OpenStudy (anonymous):

The natural log will give you the power you have to raise e to in order to get a particular value, so if you raise e to the power that you have to raise e to to get \(5x+5\), it's going to equal \(5x+5\), i.e. \(e\) cancels out \(\ln\).

OpenStudy (anonymous):

So from the above equation, you can turn it into \[5x+5 = e^2\]

OpenStudy (anonymous):

so then do you devide everything by 5? @Richard_Feynman

OpenStudy (anonymous):

Most people would subtract 5 before dividing but you can do it either way.

OpenStudy (anonymous):

You should end up with \(x = \frac{e^2}{5} - 1\).